# Multiway Alignment of Political Attitudes

### v1

@@ 5 Discussion
We build upon existing approaches for the study of pairwise alignment and provide a quantitative method to analyze alignment across multiple topics. Our proposed approach retains similar interpretability as its corresponding pairwise version, enabling the investigation of issue alignment and partisan alignment across various systems. We demonstrated its utility in contexts where opinion groups are straightforwardly defined by pre-existing categories, such as surveys. We also illustrated its applicability in instances where individual group membership is not explicitly available, such as when looking at opinion groups observed in online social networks. Additionally, when applied to Parliamentary data, our multiway alignment measure identifies subsets of issues that exhibit high alignment based on co-voting patterns and irrespective of party affiliation.
Finally, we note that the sample size required for the analysis of many topics can be a limiting factor, as multiway analysis becomes more meaningful when the size of the sample of individuals substantially exceeds the product (across topics) of the number of opinions that can be manifested at each topic.
Our work highlights how multiway alignment provides valuable insights into the dynamics of socio-political systems, revealing the presence of higher-order constraints across various contexts. While our case studies represent only a subset of the potential applications of our method, they demonstrate its accuracy and robustness. Looking ahead, our work sets the stage for future research on societal polarization and the emergence of partisan animosity. Additionally, our research equips researchers with computational tools, made available through the open-source Python package multiway_alignment, to quantify and visualize multiway alignment as we have in this paper.
@@ Appendix A Data
We operationalize the suggested multiway alignment framework across diverse datasets, including survey data (described in Appendix A.1), retweet networks (Appendix A.2), and Parliamentary roll-call votes (Appendix A.3).
@@ A.2 Retweet Networks
In line with the methodology employed in (Chen et al., 2021), we apply our algorithm to Twitter data related to Finnish political elections in 2019. Additionally, we extend our analysis to include data from the Finnish political elections in 2023 (Salloum et al., 2024).
For the 2019 dataset, tweets and retweets were gathered during the same period (March to July 2019), using a predefined set of Finnish hashtags as in (Chen et al., 2021). The collected data were then categorized into 11 distinct labels based on their text content. These labels encompass conversation streams related to Finnish political parties (SDP, National Coalition, Finns Party, Green Party, Left Party, Center Party) and salient topics (economic policy, social security, immigration, climate, education). Given that a single tweet can relate to multiple categories simultaneously, this approach results in a multilabel classification. Subsequently, for each label, we constructed the corresponding retweet network as detailed in Appendix B.2.
@@ B.1 Parliamentary Roll-Calls
For each issue or policy domain, each of the Parliament members (MPs) is represented with the vector of votes they have casted on all the roll-calls regarding the specific issue. Specifically, for each issue, each MP is associated to a vector $u_{i}=(u_{i1},u_{i2},...)$, where the element $u_{ij}$ is the vote casted by the MP $i$ on the $j$-th roll-call regarding that issue, and can take values $-1$ (no), $1$ (yes), or 00 (other, as described in Appendix A.3). This representation allows us to detect MPs with highly co-voting behavior from the cosine similarity $u_{m}.u_{n}$ between their vectors. In fact, Parliament members who cast the same votes have small cosine distance and high cosine similarity. As a result, for each issue, we can build the similarity network of the members of Parliament, where the interaction between each pair of MPs is modeled by a link weighted proportional to the similarity of their voting behavior. More precisely, given one policy domain or issue, the network of Parliament members is defined as follows:
@@ B.2 Retweet Networks
In each of the retweet networks, nodes represent users who shared a tweet (original or retweet) containing at least one relevant hashtag. Nodes are connected by an undirected edge if one user has retweeted the other on a given topic. Following the approach established in (Chen et al., 2021) and (Garimella et al., 2018), we employed the METIS algorithm (Karypis and Kumar, 1998) to separately partition each networks into two similarly-sized groups with minimal connections between them. To maintain a realistic representation of group distribution in the network, a maximum imbalance constraint of 3:7 was applied, preventing the algorithm from producing a trivial partition. Partitions for each of the retweet networks are available in Zenodo (Iannucci et al., 2024).
@@ Appendix C Maximal Alignment Curve
Appendix D Interpretation of $A_{MI}$ Measure
@@ Appendix F Null Model
$A_{net}=\frac{A-\langle A_{\text{null}}\rangle}{1-\langle A_{\text{null}}\rangle}$ (5)
@@
Appendix G Multiway Alignment of Elites: U.S. House
@@
Appendix H Use of the multiway_alignment Python Package
The multiway alignment Python package developed in conjunction with this research is available as open-source in GitHub. The code has $90\%$ coverage from unit tests and it is tested on Python $3.10$, $3.11$, and $3.12$. The multiway alignment package can be utilized in any Python environment using Python $\geq 3.10$ by following the steps outlined below. The package can be installed directly from the Python Package Index (PyPI) using pip from the command-line interface, by executing the following command:
@@ Appendix H Use of the multiway_alignment Python Package
Appendix I Alternative Formulation Using Full Consensus Partitions
@@
I.1 Expected Alignment in Case of Equal-Sized Opinion Groups
As shown above,
$A^{*}=\frac{2}{k}\sum_{i=1}^{k}\frac{1}{1+\frac{H(\text{C})}{H(T_{i})}},$
where the entropy $H=-\Sigma_{x}p_{x}\log{p_{x}}$ of a topic is 0 if all individuals belong to the same opinion group $x$. On the other hand, the highest entropy is reached when all the opinion groups are equal-sized, as in that case the probability of belonging to one group is simply $\frac{1}{G}$, where $G$ is the number of opinion groups.
Assume a social system of $N$ individuals, and $k$ topical issues, each one allowing for the same number of equally-likely opinion groups $G$. Then, the expected $k$-wise alignment can be derived in closed form and simplifies as:
$\mathbb{E}[A^{*}]=\frac{2}{k}\sum_{i=1}^{k}\frac{1}{1+\frac{\mathbb{E}[H(\text{C})]}{H(T_{i})}}$
$=\frac{2}{k}\sum_{i=1}^{k}\frac{1}{1+\frac{\mathbb{E}[H(\text{C})]}{G(-\frac{1}{G}\log{\frac{1}{G}})}}$
$=\frac{2}{1+k}.$

### v2

@@ 6 Discussion
In this paper, we introduce a quantitative method for analyzing political alignment across multiple topics, building upon existing approaches that focus on pairwise alignment. Our multiway alignment measure retains similar interpretability as traditional pairwise approaches while offering a more comprehensive view of alignment patterns in various socio-political contexts. Our method proves especially useful for investigating both issue and partisan alignment across different systems, providing insights that go beyond what can be captured with pairwise methods alone.
Our approach can be applied in diverse contexts, including public opinion surveys, Parliamentary voting records, and online discussions. Our case study on ANES data demonstrates how our measure can be applied in contexts where opinion groups are straightforwardly defined by pre-existing categories, such as surveys responses. By applying our measure, we were able to detect and quantify the emergence of well-known partisan divides. In addition to survey data, our measure can also be applied to Parliamentary voting, where we can identify subsets of issues that exhibit high alignment based on co-voting patterns and irrespective of party affiliation. Another application of our measure is the analysis of social media discussions, where individual group membership is not explicitly available but can be inferred via clustering or community detection methods.
These use cases highlight the ability of our multiway alignment measure to uncover patterns of higher-order alignment that might not visible in pairwise analyses. We have highlighted this by presenting three archetypical shapes of the alignment spectrum, which visually demonstrate how positions on one topic can inform or constrain opinions on others. These patterns help clarify how ideological and partisan divisions manifest across multiple issues, supporting the study of alignment.
The implications of this work extend both theoretically and practically. Theoretically, the multiway alignment measure enhances our ability to study political polarization, not just at the level of individual issues, but with respect to the belief systems. This is particularly relevant in understanding which issues or topics have the potential to drive polarization and societal divides. Furthermore, this approach enables research related to the complexity of ideological landscapes, allowing for more systematic assessments of how political identities and belief systems coalesce or fragment over time.
Practically, our approach equips researchers with a versatile tool for studying alignment across a range of contexts. However, some challenges remain. One limitation is the need for sufficiently large sample sizes, especially when analyzing many topics simultaneously. When the sample size is small relative to the number of topics, the results may be less robust. In such cases, we recommend comparing the alignment scores with those obtained from the null model described in Appendix F. Another challenge is the assumption that the data used to compute alignment is representative of broader societal attitudes, which may not always be the case, especially when relying on self-selecting platforms like online social media. Additionally, the computational complexity of calculating multiway alignment increases as the number of issues grows, which could pose practical constraints in large-scale datasets.
Despite these limitations, our method paves the way for several promising directions in future research. First, the multiway alignment measure could be applied in a comparative context, allowing researchers to explore whether the patterns observed in one political system hold across different systems or countries. Moreover, our measure can be used to study the evolution of alignment over time, particularly in response to major events like elections, social movements, or crises such as the COVID-19 pandemic.
In conclusion, our multiway alignment measure provides a powerful tool for understanding the structure of political attitudes across multiple dimensions. By uncovering higher-order alignment patterns, it offers new insights into the dynamics of ideological coherence, partisan divides, and issue-based polarization. While our case studies demonstrate the utility of the measure in specific contexts, its potential applications extend beyond the examples provided in this paper. We hope that this work sets the stage for future research on societal polarization and partisan animosity. To facilitate this, we offer an open-source Python package, multiway_alignment, enabling researchers across disciplines to adopt and extend this methodology.
@@ Appendix A Data
We operationalize the suggested multiway alignment framework across diverse datasets, including survey data, retweet networks, and Parliamentary roll-call votes.
@@ A.2 Retweet Networks
In line with the methodology employed in Chen et al. (2021), we apply our algorithm to Twitter data related to Finnish political elections in 2019. Additionally, we extend our analysis to include data from the Finnish political elections in 2023 (Salloum et al., 2024).
For the 2019 dataset, tweets and retweets were gathered during the same period (March to July 2019), using a predefined set of Finnish hashtags as in Chen et al. (2021). The collected data were then categorized into 11 distinct labels based on their text content. These labels encompass conversation streams related to Finnish political parties (SDP, National Coalition, Finns Party, Green Party, Left Party, Center Party) and salient topics (economic policy, social security, immigration, climate, education). Given that a single tweet can relate to multiple categories simultaneously, this approach results in a multilabel classification. Subsequently, for each label, we constructed the corresponding retweet network as detailed in B.2.
@@ B.1 Parliamentary Roll-Calls
For each issue or policy domain, each of the Parliament members (MPs) is represented with the vector of votes they have casted on all the roll-calls regarding the specific issue. Specifically, for each issue, each MP is associated to a vector $u_{i}=(u_{i1},u_{i2},...)$, where the element $u_{ij}$ is the vote casted by the MP $i$ on the $j$-th roll-call regarding that issue, and can take values $-1$ (no), $1$ (yes), or 00 (other, as described in A.3). This representation allows us to detect MPs with highly co-voting behavior from the cosine similarity $u_{m}.u_{n}$ between their vectors. In fact, Parliament members who cast the same votes have small cosine distance and high cosine similarity. As a result, for each issue, we can build the similarity network of the members of Parliament, where the interaction between each pair of MPs is modeled by a link weighted proportional to the similarity of their voting behavior. More precisely, given one policy domain or issue, the network of Parliament members is defined as follows:
@@ B.2 Retweet Networks
In each of the retweet networks, nodes represent users who shared a tweet (original or retweet) containing at least one relevant hashtag. Nodes are connected by an undirected edge if one user has retweeted the other on a given topic. Following the approach established in Chen et al. (2021) and Garimella et al. (2018), we employed the METIS algorithm (Karypis and Kumar, 1998) to separately partition each networks into two similarly-sized groups with minimal connections between them. To maintain a realistic representation of group distribution in the network, a maximum imbalance constraint of 3:7 was applied, preventing the algorithm from producing a trivial partition. Partitions for each of the retweet networks are available in Zenodo (Iannucci et al., 2024).
@@ Appendix C Maximal Alignment Curve
Appendix D Interpretation of $A_{\text{MI}}$ Measure
@@ Appendix F Null Model
$A_{\text{net}}=\frac{A-\langle A_{\text{null}}\rangle}{1-\langle A_{\text{null}}\rangle}$ (5)
@@
Appendix G Multiway Alignment in Finnish Twitter Before Elections
Appendix H Multiway Alignment of Elites: U.S. House
@@ Appendix G Multiway Alignment in Finnish Twitter Before Elections
Appendix I Multiway Alignment in Finnish Parliament Before and During COVID Pandemic
The analysis of multiway alignment within the Finnish Parliament offers insights into the shifting political dynamics before and during the COVID-19 pandemic, and reveals a known phenomenon that a purely pairwise analysis would miss. The line chart in Figure 9 shows the evolution over time of the overall alignment, summarized through the area under the maximal alignment curve. The two points highlighted in the line chart show how the emergence of the pandemic and the subsequent declaration of a state of emergency led to a sharp decrease in alignment in 2020–2021.
A comparison of the multiway alignment spectrum for the 2019–2020 and 2020–2021 parliamentary periods is shown in the main panel in Figure 9. Before the onset of the pandemic, the spectrum of the multiway alignment within the Finnish Parliament displayed a funnel shape that we commonly observe in legislative systems, where the alignment across different topics reflects the dynamics of policymaking and political discourse. The funnel shape of multiway alignment indicates a gradual decrease in alignment as the number of topics in the analysis increases. Nevertheless, the values can remain at values above zero also at higher orders (e.g., $k=9$ and $k=10$ in 2019–2020), demonstrating alignment that could originate from adherence to party lines, whereby members from the same party (or coalition) vote in a similar way on a wide range of issues.
The first hint of a reorientation of priorities and focal points within parliamentary deliberations in 2020–2021 is a shift in the most aligned pair of topics, from defense and health in 2019–-2020 to transportation and health in 2020-–2021. However, the manifest shift in the distribution of the scores already at $k=3$ usher in a period of exceptional policy discussions. At $k=3$, the spectrum for 2020–2021 shows two clear clusters of topics, one retaining alignment levels comparable to the previous period, and the other exhibiting markedly low alignment. The same clustering effect becomes increasingly distinct at higher orders ($k=4,5,6$): in 2020–2021, six topics – transportation, culture, defense, economics, foreign affairs, and health – clearly follow similar alignment patterns as in 2019-2020, while the remaining five topics – climate, education, immigration, social issues, and unemployment – become much less aligned, bringing the multiway alignment score at $k=10$ and $k=11$ towards zero.
On one hand, the sudden prominence of topics related to transportation and their alignment with health and economics suggests a focused effort within the Finnish Parliament to navigate the circumstances imposed by the pandemic through initiatives related to free movement restrictions, economic stimulus packages, quarantine and isolation protocols, and wide vaccination efforts (YLE, 2020a, b). On the other hand, topics such as climate and immigration could have been perceived as less immediately relevant or lower in priority compared to the urgent challenges posed by the crisis. Overall, the observation of the double-triangle pattern of alignment in the analysis of Finnish Parliament voting supports the claim that, throughout the pandemic, parliamentary discussions diverged from traditional political dynamics, and that the opposition adopted a collaborative approach, refraining from strong criticism of COVID policies proposed by the government (Niemikari and Raunio, 2022).
Appendix J Use of the multiway_alignment Python Package
The multiway_alignment Python package developed in conjunction with this research is available as open-source in GitHub. The code has $90\%$ coverage from unit tests and it is tested on Python $3.10$, $3.11$, and $3.12$. The multiway_alignment package can be utilized in any Python environment using Python $\geq 3.10$ by following the steps outlined below. The package can be installed directly from the Python Package Index (PyPI) using pip from the command-line interface, by executing the following command:
@@ Appendix I Multiway Alignment in Finnish Parliament Before and During COVID Pandemic
Appendix K Alternative Formulation Using Full Consensus Partitions